Time Warner could offer the History Channel (H) and Showtime (S) individually or as a bundle of both.

Suppose the reservation prices of customers 1 and 2 (the highest prices they are willing to pay) are presented in the boxes below.

The cost to Time Warner is $1 per customer for licensing fees.

Preferences

Showtime

History Chanel

Customer 1

9

2

Customer 2

3

8

Should Time Warner bundle or sell separately? Your answer needs to include the unbundled and bundled profits.
Answer #1

Solution:

Here the willingness to pay for two different customers are given for different channels. Now, if the price of “Showtime” “History Chanel” are “9” and “8” respectively, => “customer 1” will purchase only “showtime” and “customer 2” will purchase “history channel”, => the profit is given by “9+8-2 = 15”.

Now, given the information “customer 1’s” willingness to pay for the bundle of two channels is “9+2=11”. Similarly, “customer 2’s” willingness to pay for the bundle of two channels is “8+3=11”. So, if the price of bundle is set at “11”, => then the profit is given by “11*2-2=20”. So, here the profit under “bundle system” is larger compare to the profit under “single system”, => Time Warner should provide it through bundles.